OpenStudy (anonymous):

Im having another problem for my math and the question is...."Write two equivalent for 4 out of 5" but the 4 out of 5 is in fraction...so i need two equivalent ratio...

OpenStudy (aum):

If you multiply the numerator and the denominator of a fraction by the SAME number the fractions are equivalent which means the value of the fraction remains the same.

OpenStudy (aum):

\[\large \frac 45 = \frac{8}{10} = \frac{12}{15} = \frac{400}{500} = ... \]